### v1

@@ Abstract
Weakly-supervised semantic segmentation (WSSS) has achieved remarkable progress using only image-level labels. However, most existing WSSS methods focus on designing new network structures and loss functions to generate more accurate dense labels, overlooking the limitations imposed by fixed datasets, which can constrain performance improvements. We argue that more diverse trainable images provides WSSS richer information and help model understand more comprehensive semantic pattern. Therefore in this paper, we introduce a novel approach called Image Augmentation Agent (IAA) which shows that it is possible to enhance WSSS from data generation perspective. IAA mainly design an augmentation agent that leverages large language models (LLMs) and diffusion models to automatically generate additional images for WSSS. In practice, to address the instability in prompt generation by LLMs, we develop a prompt self-refinement mechanism. It allow LLMs to re-evaluate the rationality of generated prompts to produce more coherent prompts. Additionally, we insert an online filter into diffusion generation process to dynamically ensure the quality and balance of generated images. Experimental results show that our method significantly surpasses state-of-the-art WSSS approaches on the PASCAL VOC 2012 and MS COCO 2014 datasets.
@@
Weakly Supervised Learning Semantic Segmentation Diffusion Model Large Language Model
@@ 1 Introduction
Weakly supervised semantic segmentation (WSSS) leverages image-level labels to achieve dense pixels segmentation. Current mainstream WSSS methods focus on designing new network architectures and loss functions to get better Class Activation Map (CAM)[24, 35, 30]. For example, MCTformer[31] designed a transformer with multi-class tokens to generate class-specific attention maps to generate refine CAM. And the method presented in [24] incorporates a token contrastive loss for intra-class compactness and inter-class separability to mitigate over-smoothing CAM. However, these previous methods mentioned above are all only use limited scale of the available training data which restrict the performance improvement. Although previous work IACD [29] had explored using diffusion to generate augmented images for WSSS as shown in Fig. 1(b), the use of single background prompt resulted in insufficient image diversity, and the ex-post filtering operation for all generated data can easily lead to uneven number of the selected augmented images.
Figure 1: (a) In the traditional WSSS framework, the original image is directly input into WSSS. (b) In our IAA, we utilize an augmentation agent to generate additional images and then combined with the original image and input into WSSS.
Building on the aforementioned problems, we propose a novel Image Augmentation Agent (IAA) to generate diverse and high-quality even synthetic training data for WSSS as shown in Fig. 1(b). Our main contributions can be summarized as follows:
@@
    We introduce a novel augmentation agent for WSSS, leveraging the capabilities of GPT and diffusion models to generate supplementary training images.
@@
    We present a self-refinement mechanism within the agent to ensure image quality, including the dynamic refinement of background prompts and the enhancement of generated images.
@@
    Our experimental results demonstrate that this framework significantly outperforms other state-of-the-art (SOTA) methods on the PASCAL VOC 2012 and MS COCO 2014 for the segmentation task.
@@ 2.1 Weakly-Supervised Semantic Segmentation
Weakly Supervised Semantic Segmentation (WSSS) methods using image-level annotations commonly rely on Class Activation Maps (CAMs) as pseudo labels. However, CAMs often highlight only the most discriminative regions of objects, leaving less salient regions unutilized. Various approaches have been proposed to overcome this limitation, including region erasure [28], accumulating attention online [12], and mining cross-image semantics [27]. Techniques such as leveraging saliency maps [15] aim to reduce background interference and discover less obvious object regions. Additionally, contrastive methods [5] attempt to activate entire object regions by comparing pixel and prototype representations. Some studies, like [3], enhance WSSS by incorporating more category-specific information or leveraging additional learning signals from the training data. Recent advancements have explored integrating Vision Transformers (ViTs) into WSSS. For instance, MCTformer [31] utilizes ViT attention maps to create localization maps, while AFA [23] leverages multi-head self-attention and affinity modules for propagating pseudo labels. ViT-PCM [22] pioneers CAM-independent ViT applications for WSSS. These methods primarily optimize network structures or include additional features, often constrained by dataset size. In contrast, our work focuses on data augmentation, generating additional training data to advance WSSS.
@@ 2.2 Prompt-based Language Models
Prompt-based learning enhances pre-trained language models (PLMs) by appending task-specific instructions to inputs. Early strategies employed manually crafted prompts to tackle specific tasks [36]. However, manually designed prompts are inflexible and challenging to adapt to new tasks, prompting research into automatic prompt generation [25]. Continuous prompt optimization [17] has further advanced the field, improving adaptability and effectiveness across diverse tasks. Beyond text-based applications, PLMs have recently shown potential in vision tasks; for instance,  [34] uses prompts to enhance few-shot learning. Distinctly, our approach employs PLMs to generate diverse prompts, enriching textual descriptions for WSSS. To the best of our knowledge, this is the first work that utilizes self-refinement PLMs to improve WSSS through diverse prompt generation.
2.3 Diffusion Probabilistic Model
@@
Diffusion Probabilistic Models (DPMs), introduced by [26], have seen substantial progress in image generation. Latent Diffusion Models (LDMs) [20] refine this process by performing diffusion in latent spaces [7], significantly lowering computational requirements. Text-to-image diffusion models, leveraging CLIP [19] and similar pre-trained language models, have achieved remarkable image synthesis results by transforming text into latent representations. Enhanced by methods such as Stable Diffusion [21] and ControlNet [33], DPMs now generate high-quality images with precision and minimal artifacts. Recent studies [11] demonstrate the utility of DPMs in generating supplementary training data, thereby boosting task performance. Harnessing the strengths of DPMs, we propose IAA, which integrates conditional DPMs and GPT-generated prompts for WSSS. This represents the first application of conditional diffusion models in this context.
@@ 3.1 Overall framework
As illustrated in Fig. 2, the components and process of our IAA framework are illustrated. We use ControlNet [33] as the diffusion [21] backbone and GPT-4o [1] as the LLM. We pre-train the image classifier as an image selector using ViT with the original training data. The image class label is input into the LLM module, and we design a self-refinement process to improve prompt quality. The augmentation module uses the training image and the refined prompt from the LLM as input to generate the augmented image. We integrate the classifier as a filter into the diffusion step before generating the image to ensure that each augmented image is of high quality and control the class distribution of generate images. Finally, the training image combined with the augmented image is used as the final input for the WSSS task.
@@ 3.2 Prompt Generation with Self-Refinement
The motivation for generating prompts is to leverage LLM knowledge to create diverse background prompts that guide the diffusion model in generating images of different styles. Due to the instability of LLMs [18], we designed a self-refining prompt method for background prompt generation, as shown in Algorithm 1, which is the module $t(\cdot)$ in Fig. 2 . This method ensures that the generated background prompts are both reasonable and aligned with the current category through self-refinement within our agent.
@@ 5       $y_{0}=\text{GPT-4o}~{}(Initial\_prompt(p_{\text{gen}},Y))$;
In the WSSS setting, we denote training images as $X_{in}$ and their corresponding labels as $Y$. As depicted in Fig. 2, for each of the $N$ categories involved in the dataset, a pre-defined template $P_{gen}$ is served as language commands for GPT-4o [1] to generate intial text output $y_{0}$, which ensures the relevance and variety of language commands for generating background prompts.
@@
Where $p_{gen}$, as shown in Fig.3(a), is the initial text prompt for generating synthetic samples of category $Y$, and it will be self-refined in our Algorithm 1. We use the refined prompt template $p_{\text{refine}}$ in Fig.3b to assess the background prompt quality score. The output $score_{y_{0}}$, as shown in Fig.3(c), provides a score for the background. We then select the high-scoring prompt as our final output $P$, which will be used as the prompt in the diffusion models.
@@ 3.3 Diffusion with Image Self-Refinement
The motivation for using diffusion with image self-refinement is to leverage the ability of controlled diffusion to generate new images similar to the original ones, and to ensure image quality through our designed image self-refinement. By enriching the training data with additional enhanced images, we aim to improve the final performance of the WSSS task. To assess image quality, we trained a classifier using the original training data with image labels. As shown in Fig. 2, a ViT-based patch-driven classifier is first trained using the original dataset with image-level labels. To train the classifier, the input image $X_{in}$ is divided into $s$ input patches $X_{patch}\in\mathbb{R}^{d\times d\times 3}$ with a fixed size, where $s=\frac{hw}{d^{2}}$. The patch embeddings $F\in\mathbb{R}^{s\times e}$ are then obtained using a ViT encoder. Next, a weight matrix $W\in\mathbb{R}^{e\times|\mathcal{C}|}$ and a softmax function are applied to produce the prediction scores $Z\in\mathbb{R}^{s\times|\mathcal{C}|}$ for each patch:
@@
where $\mathcal{C}$ is the set of categories in the dataset. Global maximum pooling (GMP) is then used to select the highest prediction scores $\hat{y}\in\mathbb{R}^{1\times|\mathcal{C}|}$ in each class among all patches. Finally, $\hat{y}$ is utilized as the prediction scores for the image-level classification and the classifier is trained via using the multi-label classification prediction error (MCE):
$\mathcal{L}_{MCE}$ $=\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}{BCE(y_{c},\hat{y}_{c})}$ (3)
$=-\frac{1}{\mathcal{|C|}}\sum_{c\in\mathcal{C}}{y_{c}\log(\hat{y}_{c})+(1-y_{c})\log(1-\hat{y}_{c})},$
where, $\hat{y}_{c}$ is the prediction score of class $c$ and $y_{c}$ is the ground-truth label. Once the classifier finishes the training, we can use it to select the high-quality generated training data.
Next, we integrate the pre-train classifier into the image self-refinement module. The classifier is incorporated into the diffusion generation step. If the augmented images do not meet the quality criteria, we continue generating images until the desired quality is achieved, rather than applying a filter after all images have been generated. This approach ensures that the augmented images are evenly distributed across all images, preventing the randomness of diffusion from causing some images to have more augmented versions than others. As shown in Fig. 2, in the diffusion module, we utilize Stable Diffusion with ControlNet [33] as our generative model. In the data augmentation stage, an input image $X_{in}\in\mathbb{R}^{h\times w\times 3}$, a text prompt $P$ generated by the GPT self-refinement module, and a detection map $M$ are feed into diffusion $\delta(\cdot)$ to generate a new training data $X_{df}$. The detection map is an extra condition (e.g., Canny Edge [6] and Openpose [2]) to control the generation results.

### v4

@@ Abstract
Weakly Supervised Semantic Segmentation (WSSS), which utilizes only image-level annotations, has gained considerable attention for its efficiency and reduced cost. However, most existing WSSS methods focus on designing new network structures and loss functions to generate more accurate dense labels, overlooking the limitations imposed by fixed datasets, which can constrain performance improvements. We argue that more diverse trainable images provides WSSS richer information and help model understand more comprehensive semantic pattern. Therefore in this paper, we introduce a novel approach called Image Augmentation Agent (IAA) which shows that it is possible to enhance WSSS from data generation perspective. IAA mainly design an augmentation agent that leverages large language models (LLMs) and diffusion models to automatically generate additional images for WSSS. In practice, to address the instability in prompt generation by LLMs, we develop a prompt self-refinement mechanism. It allow LLMs to re-evaluate the rationality of generated prompts to produce more coherent prompts. Additionally, we insert an online filter into diffusion generation process to dynamically ensure the quality and balance of generated images. Experimental results show that our method significantly surpasses state-of-the-art WSSS approaches on the PASCAL VOC 2012 and MS COCO 2014 datasets. Our source code will be released.
@@
Weakly-Supervised Learning, Semantic Segmentation, Large Language Model , Diffusion Model
@@ 1 Introduction
WSSS leverages image-level labels to perform dense pixel-wise segmentation, making it a cost-effective alternative to fully supervised methods, which often require expensive pixel-wise annotations. The key advantage of WSSS lies in its ability to train segmentation models with minimal supervision, relying solely on image-level labels that provide less granular information but can still guide the model towards accurate segmentation. This approach has gained significant attention in recent years, as it allows for scaling segmentation tasks to large datasets where pixel-wise annotations are unavailable or impractical to obtain. Current mainstream WSSS methods primarily focus on improving the generation of Class Activation Maps (CAMs), which serve as weak supervisory signals for segmentation. These methods typically involve designing novel network architectures and loss functions that enhance the quality and effectiveness of CAMs, which in turn improves segmentation accuracy Wu et al. [2025a], Ru et al. [2023], Zhao et al. [2024], Yin et al. [2023], Wu et al. [2024b]. For example, MCTformer Xu et al. [2022] introduced a transformer-based architecture with multi-class tokens to generate class-specific attention maps. This enables a more refined representation of the image, allowing for better localization of objects and more precise CAMs. Similarly, the method presented in Ru et al. [2023] incorporates a token contrastive loss, which enhances intra-class compactness and inter-class separability. This approach mitigates the problem of over-smoothing in CAM generation, where objects of different classes can be confused with one another due to the lack of fine-grained supervision. Despite these advancements, WSSS continues to face two major challenges: (1) the lack of detailed supervision, such as pixel-level annotations, which hampers the model’s ability to accurately localize object boundaries, and (2) the limited scale and diversity of training data, which restricts generalization to unseen categories and complex scenes. These issues result in learning bottlenecks, slow performance gains, and limited robustness. To mitigate these limitations, recent efforts have explored data augmentation using synthetic image generation and external information sources. Notably, IACD Wu et al. [2024a] applied diffusion models to generate additional images for training, as shown in Fig. 1(b). However, IACD suffers from several drawbacks: it employs a fixed background prompt that limits visual diversity, and its post-hoc filtering process may cause an uneven sample distribution, leading to suboptimal dataset quality.
Figure 1: (a) In the traditional WSSS pipeline, the original image is directly used for training. (b) In our proposed IAA framework, an augmentation agent generates additional synthetic images, which are combined with the original image to enhance training for WSSS.
To address the above challenges, we propose a novel Image Augmentation Agent (IAA) that automatically generates diverse, high-quality synthetic training data for WSSS. Our agent integrates LLM-driven prompt generation with a diffusion model and introduces a self-refinement mechanism that enhances both prompt diversity and image quality. The entire pipeline is automated and designed to produce semantically rich and balanced synthetic datasets, as illustrated in Fig. 1(b).
Our main contributions are summarized as follows:
@@
    We propose a novel augmentation framework for WSSS that synergizes large language models (LLMs) and diffusion models to generate diverse and semantically consistent training images.
@@
    We design a self-refinement mechanism to enhance prompt diversity and improve image quality, addressing the limitations of fixed-prompt and post-filtering-based approaches.
@@
    Extensive experiments on PASCAL VOC 2012 and MS COCO 2014 benchmarks show that our method achieves consistent improvements over existing state-of-the-art methods.
@@ 2.1 Weakly-Supervised Semantic Segmentation
WSSS methods using image-level annotations commonly rely on CAMs as pseudo labels. However, CAMs often highlight only the most discriminative regions of objects, leaving less salient regions unutilized. Various approaches have been proposed to overcome this limitation, including region erasure Wei et al. [2017], accumulating attention online Jiang & Hou [2019], and mining cross-image semantics Sun et al. [2020]. Techniques such as leveraging saliency maps Lee et al. [2021] aim to reduce background interference and discover less obvious object regions. Additionally, contrastive methods Chen et al. [2022] attempt to activate entire object regions by comparing pixel and prototype representations. Some studies, like Chang et al. [2020], Wu et al. [2025b, 2024c], Long et al. [2025, 2024, 2023], Ge et al. [2025], Zhang et al. [2024], enhance WSSS by incorporating more category-specific information or leveraging additional learning signals from the training data. Recent advancements have explored integrating Vision Transformers (ViTs) into WSSS. For instance, MCTformer Xu et al. [2022] utilizes ViT attention maps to create localization maps, while AFA Ru et al. [2022] leverages multi-head self-attention and affinity modules for propagating pseudo labels. ViT-PCM Rossetti et al. [2022] pioneers CAM-independent ViT applications for WSSS. WeakCLIP Zhu et al. [2025a] introduces a novel CLIP-based text-to-pixel matching framework for WSSS, leveraging vision-language pretraining to refine CAMs and improve segmentation performance. Moreover, recent transformer-based methods have addressed over-smoothing in ViTs to improve segmentation quality. For example, Cheng et al. Cheng et al. [2025] propose a calibration framework to mitigate over-smoothing and enhance discriminative power in WSSS, achieving strong performance on benchmark datasets.
These methods primarily optimize network structures or include additional features. In contrast, our work takes a data-centric perspective by focusing on data augmentation through synthetic generation. Rather than modifying model architectures, we aim to enhance WSSS performance by expanding the training corpus in a structured way. Our approach is complementary to existing architectural improvements and can serve as a plug-and-play module compatible with various WSSS pipelines, including transformer-based ones.
@@ 2.2 Prompt-based Language Models
Prompt-based learning enhances pre-trained language models (PLMs) by appending task-specific instructions to inputs, allowing the model to better adapt to a wide range of tasks. Early strategies primarily focused on manually crafting prompts that were designed to address specific tasks Zou et al. [2021], Zhu et al. [2024, 2025b], Li et al. [2024], Guo et al. [2024]. These manually designed prompts proved effective in certain domains but were inherently limited by their lack of flexibility. As a result, they were difficult to generalize across new or unseen tasks. This limitation spurred research into automating the generation of prompts, allowing for more scalable and adaptable solutions Shin et al. [2020], Guo et al. [2025]. Through the development of automatic prompt generation methods, it became possible to generate prompts dynamically based on the task at hand, enhancing the model’s ability to generalize and improving its performance across a variety of domains. One significant advancement in the field was the introduction of continuous prompt optimization Liu et al. [2023], which further enhanced the adaptability and effectiveness of prompts. By optimizing prompts in a continuous space rather than a discrete one, models can now generate more precise and task-relevant prompts, allowing for greater flexibility in handling a diverse range of tasks. This technique has proven to be effective in improving performance in natural language processing (NLP) tasks, and its principles have since been applied to other areas of machine learning. In addition to their success in text-based applications, PLMs have recently demonstrated significant potential in vision-related tasks. A notable example of this is the use of prompts to enhance few-shot learning for visual recognition Zhang et al. [2023b]. By generating task-specific textual prompts that guide the model in interpreting visual input, PLMs have opened new possibilities for cross-domain applications. Unlike traditional methods that rely heavily on manually designed features, this approach leverages the inherent power of language models to understand and generate context-specific instructions that improve task performance. Distinctly, our approach introduces a novel use of PLMs to generate diverse prompts specifically for enhancing WSSS. By generating a variety of prompts that enrich the textual descriptions associated with images, our method aims to improve the performance of WSSS in a way that was not previously explored. To the best of our knowledge, this is the first work to apply self-refinement techniques in PLMs to generate diverse prompts, thereby improving WSSS tasks. This approach not only enriches the textual descriptions used for model training but also contributes to a more effective use of weakly labeled data for segmentation purposes.
2.3 Diffusion Models for Data-Efficient Learning
@@
Diffusion Probabilistic Models (DPMs), introduced by Sohl-Dickstein et al. [2015], have seen substantial progress in image generation. Latent Diffusion Models (LDMs) Richardson et al. [2021] refine this process by performing diffusion in latent spaces Esser et al. [2021], significantly lowering computational requirements. Text-to-image diffusion models, leveraging CLIP Radford et al. [2021] and similar pre-trained language models, have achieved remarkable image synthesis results by transforming text into latent representations. Enhanced by methods such as Stable Diffusion Rombach et al. [2022] and ControlNet Zhang & Agrawala [2023], DPMs now generate high-quality images with precision and minimal artifacts.
Recent studies Ho et al. [2020] demonstrate the utility of DPMs in generating supplementary training data, thereby boosting task performance. In parallel to diffusion advancements, recent vision research increasingly explores prompt-driven strategies for improving performance in low-data settings. For instance, M-RRFS Huang et al. [2024] introduces a memory-based region feature synthesizer for zero-shot object detection, addressing intra- and inter-class variations through prompt-inspired structures. Similarly, UPLVP Zhang et al. [2025] presents an unsupervised pre-training framework using language-vision prompts to enhance instance segmentation with limited supervision.
These works highlight the emerging trend of combining prompts and generative models (e.g., diffusion) for data-efficient learning across vision tasks. Motivated by this direction, we propose IAA, which integrates conditional DPMs with GPT-generated prompts to generate high-quality and diverse training images for WSSS. To the best of our knowledge, this represents the first application of conditional diffusion models for WSSS.
@@ 3.1 Overall framework
As illustrated in Fig. 2, the components and process of our IAA framework are illustrated. We use ControlNet Zhang & Agrawala [2023] as the diffusion Rombach et al. [2022] backbone and GPT-4o Hurst et al. [2024] as the LLM. We pre-train the image classifier as an image selector using ViT with the original training data. The image class label is input into the LLM module, and we design a self-refinement process to improve prompt quality. The augmentation module uses the training image and the refined prompt from the LLM as input to generate the augmented image. We integrate the classifier as a filter into the diffusion step before generating the image to ensure that each augmented image is of high quality and control the class distribution of generate images. Finally, the training image combined with the augmented image is used as the final input for the WSSS task.
@@ 3.2 Prompt Generation with Self-Refinement
The motivation behind our prompt generation strategy lies in fully harnessing the extensive world knowledge embedded within LLMs to construct a wide range of background prompts. These prompts serve as critical guides for the diffusion model, enabling it to generate images with diverse styles, semantics, and visual properties. By relying on the generative capabilities of LLMs, we aim to inject rich variability into the prompt space, which in turn supports the creation of more flexible and expressive image outputs. This diversity is especially crucial in vision tasks such as image generation and WSSS, where precise control over visual content is needed.
However, despite their powerful generative capacity, LLMs are not without flaws. A well-known limitation is their inherent instability Madaan et al. [2024], which can result in the generation of incoherent, overly generic, or irrelevant prompts that misalign with the intended semantic category. Such low-quality prompts can severely compromise the performance of downstream image generation, particularly when task-specific fidelity is required. To address this challenge, we designed a self-refining prompt method that refines the generated prompts iteratively, ensuring that the background prompts are not only reasonable but also closely aligned with the specific category they correspond to. This iterative refinement process allows for the correction of inconsistencies or errors in the initial prompts generated by the LLM, thereby improving their relevance and quality. As shown in Algorithm 1, this self-refining method plays a central role in the prompt generation pipeline. It is represented as the module $t(\cdot)$ in Fig. 2, where it interacts with the LLM to refine the background prompts continuously. The self-refinement process works by taking the initial prompt generated by the LLM and evaluating its quality based on a predefined set of criteria. If the prompt does not meet the required standards, it is refined further, ensuring that it remains consistent with the category label and maintains its effectiveness in guiding the diffusion model. Through this process, the prompts are iteratively adjusted until they meet the desired level of quality, making them more effective in generating images that are diverse yet relevant to the given task.
This approach not only helps improve the quality of the generated background prompts but also increases the robustness and reliability of the entire image generation process. By ensuring that the prompts align with the current category and are continuously refined, we can guide the diffusion model more effectively and generate images with greater accuracy and diversity. This self-refining method is crucial for handling the complexities of generating high-quality images in tasks such as WSSS, where the ability to generate coherent and diverse images is paramount.
@@ 5 $y_{0}=\text{GPT-4o}~(Initial\_prompt(p_{\text{gen}},Y))$;
In the WSSS setting, we denote training images as $X_{in}$ and their corresponding labels as $Y$. As depicted in Fig. 2, for each of the $N$ categories involved in the dataset, a pre-defined template $P_{gen}$ is served as language commands for GPT-4o Hurst et al. [2024] to generate intial text output $y_{0}$, which ensures the relevance and variety of language commands for generating background prompts.
@@
Where $p_{gen}$, as shown in Fig.3(a), is the initial text prompt for generating synthetic samples of category $Y$, and it will be self-refined in our Algo. 1. We use the refined prompt template $p_{\text{refine}}$ in Fig.3b to assess the background prompt quality score. The output $score_{y_{0}}$, as shown in Fig.3(c), provides a score for the background. We then select the high-scoring prompt as our final output $P$, which will be used as the prompt in the diffusion models.
@@ 3.3 Diffusion with Image Self-Refinement
The motivation for using diffusion with image self-refinement is to leverage the ability of controlled diffusion to generate new images that are similar to the original ones while ensuring image quality through our designed image self-refinement mechanism. This combination allows us to maintain the diversity of the synthetic images while ensuring they are of high quality, which is crucial for improving the performance of WSSS. By enriching the training data with additional enhanced images that resemble real-world variations, we aim to improve the model’s ability to handle complex and varied inputs, ultimately improving the final performance of the WSSS task. The primary goal of controlled diffusion is to generate synthetic images that capture the inherent variability found in the original dataset, but with additional alterations introduced by the diffusion process. This approach not only enhances the model’s exposure to a wider variety of visual scenarios but also helps it adapt better to unseen variations, which are typical in real-world applications of WSSS. The image self-refinement step ensures that the generated images maintain high fidelity to the original content, which is crucial for ensuring that the model’s performance is not compromised by the introduction of low-quality synthetic images.
To further reduce the domain gap between synthetic and real images, we employ ControlNet to guide the diffusion process using structural conditions such as edge or pose maps derived from the original image. This ensures that the augmented images retain core structural semantics while introducing controlled variations. Although our method does not explicitly use CLIP-based filtering, the use of a structure-aware refinement loop inherently encourages semantic consistency with the original dataset, which helps minimize potential distribution shifts introduced by synthetic augmentation.
To assess the quality of the generated images, we trained a classifier using the original training data with image-level labels. As shown in Fig. 2, we use a Vision Transformer (ViT)-based patch-driven classifier to perform the evaluation. The classifier is first trained using the original dataset, which contains images labeled at the image level. The input image $X_{in}$ is divided into $s$ input patches $X_{patch}\in\mathbb{R}^{d\times d\times 3}$ with a fixed size, where $s=\frac{hw}{d^{2}}$ and $h$ and $w$ represent the height and width of the image. The goal is to extract meaningful patch embeddings that capture the local features of the image.
The patch embeddings $F\in\mathbb{R}^{s\times e}$ are then computed using a ViT encoder, which processes the image patches and converts them into high-dimensional representations. A weight matrix $W\in\mathbb{R}^{e\times|\mathcal{C}|}$ is used to map the embeddings into the class space, where $\mathcal{C}$ is the set of categories in the dataset. This weight matrix is applied alongside a softmax function to produce the prediction scores $Z\in\mathbb{R}^{s\times|\mathcal{C}|}$ for each patch:
@@
where $\mathcal{C}$ is the set of categories in the dataset. The softmax function normalizes the scores, ensuring they represent class probabilities for each patch. Subsequently, global maximum pooling (GMP) is applied to each class to select the highest prediction scores $\hat{y}\in\mathbb{R}^{1\times|\mathcal{C}|}$ among all the patches. This pooling operation helps summarize the class-wise information and produces a single set of image-level prediction scores.
Finally, the vector $\hat{y}$, which contains the image-level prediction scores, is used for image-level classification. The classifier is trained using the multi-label classification prediction error (MCE) loss function. This loss function evaluates the difference between the predicted scores and the ground truth labels for each image. The MCE loss is defined as:
$\mathcal{L}_{MCE}$ $=\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}\text{BCE}(y_{c},\hat{y}_{c})$ (3)
$=-\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}\left[y_{c}\log(\hat{y}_{c})+(1-y_{c})\log(1-\hat{y}_{c})\right],$
where $y_{c}$ is the ground-truth label for class $c$ and $\hat{y}_{c}$ is the predicted probability for class $c$. The binary cross-entropy (BCE) loss function is applied for each class independently. Once the classifier has been trained, it can be used to assess the quality of the generated images by evaluating how well they match the ground truth labels. The classifier’s performance in distinguishing between correct and incorrect class predictions serves as an important metric for selecting high-quality generated training data.
Using this classifier, we can filter and select the high-quality augmented images, which will then be incorporated into the final training dataset. These selected high-quality samples are combined with the original dataset to form a robust training set, which can then be used to train the WSSS model. The inclusion of high-quality augmented images significantly contributes to improving the model’s ability to perform accurate and precise semantic segmentation, especially in scenarios where labeled data is scarce or expensive to obtain.
Next, we integrate the pre-train classifier into the image self-refinement module. The classifier is incorporated into the diffusion generation step. If the augmented images do not meet the quality criteria, we continue generating images until the desired quality is achieved, rather than applying a filter after all images have been generated. This approach ensures that the augmented images are evenly distributed across all images, preventing the randomness of diffusion from causing some images to have more augmented versions than others. As shown in Fig. 2, in the diffusion module, we utilize Stable Diffusion with ControlNet Zhang & Agrawala [2023] as our generative model. In the data augmentation stage, an input image $X_{in}\in\mathbb{R}^{h\times w\times 3}$, a text prompt $P$ generated by the GPT self-refinement module, and a detection map $M$ are feed into diffusion $\delta(\cdot)$ to generate a new training data $X_{df}$. The detection map is an extra condition (e.g., Canny Edge Ding & Goshtasby [2001] and Openpose Cao et al. [2017]) to control the generation results.